OpenStudy (anonymous):

Simplify completely quantity 2 x squared plus 16 x plus 30 all over quantity 5 x squared plus 13 x minus 6

OpenStudy (stamp):

\[\frac{2x^2+16x+30}{5x^2+13x-6}\]\[\frac{(2x+6)(x+5)}{(5x-2)(x+3)}\]

OpenStudy (anonymous):

im doing a worksheet and thats not one of the options

OpenStudy (stamp):

@blondiee98 because that is not the final simplification. i am not here to do your homework for you

OpenStudy (anonymous):

im not asking for you to do im asking for help on how to do it. i see what your saying but when i go back to my work it doesnt add up

OpenStudy (stamp):

(2x+6) = 2(x+3)

OpenStudy (stamp):

though you didn't need me to tell you that

OpenStudy (stamp):

because you knew that 2 and 6 were both even meaning they share a common factor of 2

OpenStudy (anonymous):

but then how would bottom be simplified since you said it wasnt fully simplified. i dont see how

OpenStudy (stamp):

(x+3)/(x+3) = 1 but you already know that because anything divide by itself is equal to one.